# Fractional $p$-Laplace Systems with critical Hardy Nonlinearities: Existence and Multiplicity

### v1

@@ Abstract.
{$&(-\Delta_{p})^{s}u=\frac{\alpha}{q}\frac{|u|^{\alpha-2}u|v|^{\beta}}{|x|^{m}}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\frac{\beta}{q}\frac{|v|^{\beta-2}v|u|^{\alpha}}{|x|^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega,\end{aligned}\right.$
@@
{$&(-\Delta_{p})^{s}u=\eta|u|^{r-2}u+\gamma\frac{\alpha}{p_{s}^{*}(m)}\frac{|u|^{\alpha-2}u|v|^{\beta}}{|x|^{m}}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\eta|v|^{r-2}v+\gamma\frac{\beta}{p^{*}_{s}(m)}\frac{|v|^{\beta-2}v|u|^{\alpha}}{|x|^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega,\end{aligned}\right.$
@@ 1. Introduction
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set containing 00. For $0<s<1<p<\infty$ and $d>sp$, this paper deals with the following critical system of equations driven by the fractional $p$-Laplace operators:
{$&(-\Delta_{p})^{s}u=\eta\lvert u\rvert^{r-2}u+\gamma\frac{\alpha}{q}\frac{|u|^{\alpha-2}u|v|^{\beta}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\eta\lvert v\rvert^{r-2}v+\gamma\frac{\beta}{q}\frac{|v|^{\beta-2}v|u|^{\alpha}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega,\end{aligned}\right.$ ($\mathcal{S}$)
@@
$(-\Delta)_{p}^{s}u(x)=2\lim_{\varepsilon\rightarrow 0^{+}}\int_{\mathbb{R}^{d}\backslash B_{\varepsilon}(x)}\frac{|u(x)-u(y)|^{p-2}(u(x)-u(y))}{|x-y|^{d+sp}}\,\,{\rm d}y,\quad\text{for}\leavevmode\nobreak\ x\in\mathbb{R}^{d},$
@@
$[u]_{s,p}^{p}:=\iint\limits_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,\,{\rm d}x\,{\rm d}y.$
@@
The space ${\mathcal{D}}^{s,p}(\mathbb{R}^{d})$ is a reflexive Banach space under the norm $[\cdot]_{s,p}$. Now consider the following closed subspace of ${\mathcal{D}}^{s,p}(\mathbb{R}^{d})$:
@@
endowed with the norm $[\cdot]_{s,p}$, which is equivalent to the fractional Sobolev norm $\|\cdot\|_{L^{p}(\Omega)}+[\cdot]_{s,p}$. For further details on fractional Sobolev spaces and associated embedding results, we refer to [27, 16] and the references therein. We consider $\mathcal{W}=W_{0}^{s,p}(\Omega)\times W_{0}^{s,p}(\Omega)$ as the solution space for ($\mathcal{S}$), which is endowed with the following norm
@@
{$&(-\Delta_{p})^{s}u=\frac{\alpha}{q}\frac{|u|^{\alpha-2}u|v|^{\beta}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\frac{\beta}{q}\frac{|v|^{\beta-2}v|u|^{\alpha}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega.\end{aligned}\right.$ ($\mathcal{S}_{H}$)
@@
$S_{\alpha+\beta}:=\inf_{u\in{W_{0}^{s,p}(\Omega)}\setminus\{0\}}\frac{[u]_{s,p}^{p}}{\left(\int_{\Omega}\frac{\lvert u(x)\rvert^{q}}{\lvert x\rvert^{m}}\,{\rm d}x\right)^{\frac{p}{q}}}\text{ and }S_{\alpha,\beta}:=\inf_{(u,v)\in\mathcal{W}\setminus\{0,0\}}\frac{\lVert(u,v)\rVert^{p}_{{\mathcal{W}}}}{\left(\int_{\Omega}\frac{\lvert u\rvert^{\alpha}\lvert v\rvert^{\beta}}{\lvert x\rvert^{m}}\,{\rm d}x\right)^{\frac{p}{q}}}.$
Observe that, if $u$ minimizes $S_{\alpha+\beta}$, then so is $\lvert u\rvert$, since $[\lvert u\rvert]_{s,p}\leq[u]_{s,p}$. So, without loss of generality, we assume that a minimizer of $S_{\alpha+\beta}$ is nonnegative. A similar fact holds for a minimizer of $S_{\alpha,\beta}$. If $(u,v)\in\mathcal{W}$ minimizes $S_{\alpha,\beta}$, then it satisfies
@@
{$&(-\Delta_{p})^{s}u=\eta\lvert u\rvert^{r-2}u+\gamma\frac{\lvert u\rvert^{p^{*}_{s}(m)-2}u}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega.\end{aligned}\right.$ (1.2)
@@
{$&(-\Delta_{p})^{s}u=\eta_{1}\lvert u\rvert^{q-2}u+\frac{\alpha}{p_{s}^{*}}|u|^{\alpha-2}u|v|^{\beta}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\eta_{2}\lvert v\rvert^{q-2}v+\frac{\beta}{p^{*}_{s}}|v|^{\beta-2}v|u|^{\alpha}\;\;\text{in}\;\Omega,\\ &u=v=0\;\;\text{in}\;\mathbb{R}^{d}\setminus\Omega,\end{aligned}\right.$ (1.3)
@@
In this paper, we extend the result of [14, Theorem 1.2] in the context of the system. Specifically, we prove the existence of a nontrivial solution with the least energy for ($\mathcal{S}$). Moreover, we prove that the nontrivial solution is indeed non semi-trivial. In order to prove the non semi-triviality, we note that the approach used in [13], which relies on the presence of a concave nonlinearity, does not apply to our setting. Instead, we employ an alternative approach based on the properties of the Nehari manifold.
@@
    ($\mathcal{S}$) admits a nontrivial weak solution $(\tilde{u},\tilde{v})$ with the least energy provided
    $\begin{cases}r\geq p,&\text{ if }d\geq p^{2}s;\\ r>p^{*}-p^{\prime},&\text{ if }d<p^{2}s.\end{cases}$
@@
    Further, there exists $0<\eta^{*}<\min\{\lambda_{1},\frac{S_{\alpha+\beta}}{c_{0}}\}$ such that for every $\eta\in(0,\eta^{*})$, $(\tilde{u},\tilde{v})$ is non semitrivial, i.e., $(\tilde{u},\tilde{v})\neq(\tilde{u},0)$ or $(\tilde{u},\tilde{v})\neq(0,\tilde{v})$ a.e. in $\Omega$.
@@
$-\Delta_{p}u=\mu u^{q-1}+u^{p^{*}-1}\text{ in }\Omega,\quad 2\leq p\leq q<p^{*}=\frac{dp}{d-p}.$
@@
$-\Delta_{p}u+(-\Delta_{p})^{s}u=\lambda|u|^{q-2}u+|u|^{p^{*}-2}u\text{ in }\Omega,\;u=0\text{ in }\mathbb{R}^{d}\setminus\Omega,\quad 2\leq p<q<p^{*}=\frac{dp}{d-p},$
@@
Let $\Omega$ be a bounded open set containing zero. Let $2\leq p\leq r<p_{s}^{*}$ and $\alpha,\beta\geq 2$ be such that $\alpha+\beta=p^{*}_{s}$. Let $\gamma>0$ and $r$ be as in Theorem 1.4. Then there exists $\eta_{*}>0$ satisfying
@@
    For $a\in(1,\infty)$, $a^{\prime}:=\frac{a}{a-1}$ is the conjugate exponent of $a$.
@@ 2. Existence and Non-existence of ground state solutions
$\int_{\Omega}\frac{\lvert u(x)\rvert^{\alpha}\lvert v(x)\rvert^{\beta}}{\lvert x\rvert^{m}}\,{\rm d}x=\int_{\Omega}\frac{\lvert u(x)\rvert^{\alpha}}{\lvert x\rvert^{\frac{m\alpha}{q}}}\frac{\lvert v(x)\rvert^{\alpha}}{\lvert x\rvert^{\frac{m\beta}{q}}}\,{\rm d}x$
@@
Let $0\leq m\leq sp$, $p\leq q<p^{*}_{s}(m)$ and $\alpha,\beta>1$ be such that $\alpha+\beta=q$. Let $\{(u_{n},v_{n})\}$ weakly converges to ${(u,v)}$ in $\mathcal{W}$. Then
@@
    Let =${\mathcal{M}}=\left\{(u,v)\in{\mathcal{W}}:G(u,v)=1\right\}$. Then $S_{\alpha,\beta}=\inf\left\{\lVert(u,v)\rVert^{p}_{{\mathcal{W}}}:(u,v)\in{\mathcal{M}}\right\}$ admits a positive minimizer in ${\mathcal{M}}$.
@@
Similarly, we also get $\int_{\Omega}\frac{\lvert v_{n}-v\rvert^{q}}{|x|^{m}}\,{\rm d}x\rightarrow 0$ as $n\rightarrow\infty$. Therefore, taking $n\rightarrow\infty$ in (2.7) and (2.8), we conclude that $G$ is compact.
@@
Let $0\leq m\leq sp$, $p\leq q\leq p^{*}_{s}(m)$ and $\alpha,\beta>1$ be such that $\alpha+\beta=q$. Let $w_{0}$ be a minimizer of $S_{\alpha+\beta}$. Then for every $B,C>0$ with $\frac{B}{C}=(\alpha\beta^{-1})^{\frac{1}{p}}$,
@@
for $\text{for }\tilde{t}=\left(\frac{X}{Y}\right)^{\frac{1}{q-p}},\;E(\tilde{t}u,\tilde{t}v)=\max_{t>0}E(tu,tv).$
@@
Let $\alpha+\beta=p^{*}_{s}$. Then $S_{\alpha,\beta}$ has no minimizer on any open set $\Omega\subsetneq\mathbb{R}^{d}$.
@@
Observe that $\|(\cdot,\cdot)\|_{{\mathcal{W}}}$ and $\int_{\Omega}|\cdot|^{\alpha}|\cdot|^{\beta}\,{\rm d}x$ are invariant under translation and scaling
@@
Arguing as in [32], $S_{\alpha,\beta}$ is independent of $\Omega$ in $\mathbb{R}^{d}$. Suppose $(u_{0},v_{0})$ is a minimizer of $S_{\alpha,\beta}$ on any open set $\Omega$ in $\mathbb{R}^{d}$. Then the pair of functions $(\tilde{u},\tilde{v})\in\mathcal{W}(\mathbb{R}^{d})$ with $\tilde{u}=u_{0},\tilde{v}=v_{0}$ in $\Omega$, and $\tilde{u},\tilde{v}=0$ in $\mathbb{R}^{d}\setminus\Omega$, minimizes $S_{\alpha,\beta}$ over $\mathbb{R}^{d}$. Moreover, up to the normalized pair (as in Remark 1.1), they satisfy
{$&(-\Delta_{p})^{s}\tilde{u}=\frac{\alpha}{p^{*}_{s}}|\tilde{u}|^{\alpha-2}\tilde{u}|\tilde{v}|^{\beta}\;\;\text{in}\;\mathbb{R}^{d},\\ &(-\Delta_{p})^{s}\tilde{v}=\frac{\beta}{p^{*}_{s}}|\tilde{v}|^{\beta-2}\tilde{v}|\tilde{u}|^{\alpha}\;\;\text{in}\;\mathbb{R}^{d},\\ &\tilde{u},\tilde{v}\geq 0\;\;\text{in}\;\mathbb{R}^{d}.\end{aligned}\right.$ (2.11)
@@
Let $\alpha,\beta>1$ be such that $\alpha+\beta=p^{*}_{s}$. Let $\{(u_{n},v_{n})\}$ be a sequence in ${\mathcal{W}}$ such that
@@
has a convergent subsequence, still denoted by $(\tilde{u}_{n},\tilde{v}_{n})$, such that $(\tilde{u}_{n},\tilde{v}_{n})\to(\tilde{u},\tilde{v})$ in ${\mathcal{W}}(\mathbb{R}^{d})$. Moreover, $r_{n}\to 0$ and $y_{n}\to y$ in $\overline{\Omega}$.
@@
Following the proof of [8, Theorem 1.6, pg. 25], we see that there exists $\delta_{1}>0$, such that
$\sup_{x\in{\mathbb{R}}^{d},r>0}r^{d-sp}\fint_{B(x,r)}(|{u}_{n}|^{p}+|{v}_{n}|^{p})\,{\rm d}x=\delta_{1}.$
Hence there exist $y_{n}\in{\mathbb{R}}^{d}$, $r_{n}\in{\mathbb{R}}^{+}$ such that
$r_{n}^{d-sp}\fint_{B(y_{n},r_{n})}(|{u}_{n}|^{p}+|{v}_{n}|^{p})\,{\rm d}x=\frac{\delta_{1}}{2}.$ (2.12)
Since $u_{n},v_{n}$ vanish on $\Omega^{c}$, by (2.12), we, in fact, have $y_{n}\in\{x\in{\mathbb{R}}^{d}:d(x,\overline{\Omega})\leq r_{n}\}$. We set
$(\tilde{u}_{n},\tilde{v}_{n})=r_{n}^{\frac{d-sp}{p}}({u}_{n}(r_{n}x+y_{n}),{v}_{n}(r_{n}x+y_{n})).$
Following the step-2 of [8, Theorem 1.6], we see that $(\tilde{u}_{n},\tilde{v}_{n})\rightharpoonup(\tilde{u},\tilde{v})\neq(0,0)$ in ${\mathcal{W}}$. By Brezis Lieb lemma,
@@
11 11 $=\lim_{n\to\infty}\int_{{\mathbb{R}}^{d}}|\tilde{u}_{n}|^{\alpha}|\tilde{v}_{n}|^{\beta}\,{\rm d}x=\lim_{n\to\infty}\int_{{\mathbb{R}}^{d}}|\tilde{u}_{n}-\tilde{u}|^{\alpha}|\tilde{v}_{n}-\tilde{v}|^{\beta}\,{\rm d}x+\int_{{\mathbb{R}}^{d}}|\tilde{u}|^{\alpha}|\tilde{v}|^{\beta}\,{\rm d}x$ (2.14)

### v3

@@ Abstract.
$\left\{\begin{aligned} &(-\Delta_{p})^{s}u=\frac{\alpha}{q}\frac{|u|^{\alpha-2}u|v|^{\beta}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\frac{\beta}{q}\frac{|v|^{\beta-2}v|u|^{\alpha}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega,\end{aligned}\right$.
@@
$\left\{\begin{aligned} &(-\Delta_{p})^{s}u=\eta\lvert u\rvert^{r-2}u+\gamma\frac{\alpha}{p_{s}^{*}(m)}\frac{|u|^{\alpha-2}u|v|^{\beta}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\eta\lvert v\rvert^{r-2}v+\gamma\frac{\beta}{p^{*}_{s}(m)}\frac{|v|^{\beta-2}v|u|^{\alpha}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega,\end{aligned}\right$.
@@ 1. Introduction
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set containing 0. For $0<s<1<p<\infty$ and $d>sp$, this paper deals with the following critical system of equations driven by the fractional $p$-Laplace operators:
$\left\{\begin{aligned} &(-\Delta_{p})^{s}u=\eta\lvert u\rvert^{r-2}u+\gamma\frac{\alpha}{q}\frac{|u|^{\alpha-2}u|v|^{\beta}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\eta\lvert v\rvert^{r-2}v+\gamma\frac{\beta}{q}\frac{|v|^{\beta-2}v|u|^{\alpha}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega,\end{aligned}\right$. ($\mathcal{S}$)
@@
$(-\Delta)_{p}^{s}u(x)=2\lim_{\varepsilon\rightarrow 0^{+}}\int_{\mathbb{R}^{d}\backslash B_{\varepsilon}(x)}\frac{|u(x)-u(y)|^{p-2}(u(x)-u(y))}{|x-y|^{d+sp}}\,\,{\rm d}y,\quad\text{for}\penalty 10000\ x\in\mathbb{R}^{d},$
@@
$u]_{s,p}^{p}:=\iint\limits_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,\,{\rm d}x\,{\rm d}y.$
@@
The space ${\mathcal{D}}^{s,p}(\mathbb{R}^{d})$ is a reflexive Banach space under the norm $\cdot]_{s,p}$. Now consider the following closed subspace of ${\mathcal{D}}^{s,p}(\mathbb{R}^{d})$:
@@
endowed with the norm $\cdot]_{s,p}$, which is equivalent to the fractional Sobolev norm $\|\cdot\|_{L^{p}(\Omega)}+[\cdot]_{s,p}$. For further details on fractional Sobolev spaces and associated embedding results, we refer to [27, 16] and the references therein. We consider $\mathcal{W}=W_{0}^{s,p}(\Omega)\times W_{0}^{s,p}(\Omega)$ as the solution space for ($\mathcal{S}$), which is endowed with the following norm
@@
$\left\{\begin{aligned} &(-\Delta_{p})^{s}u=\frac{\alpha}{q}\frac{|u|^{\alpha-2}u|v|^{\beta}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\frac{\beta}{q}\frac{|v|^{\beta-2}v|u|^{\alpha}}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega.\end{aligned}\right$. ($\mathcal{S}_{H}$)
@@
$S_{\alpha+\beta}:=\inf_{u\in{W_{0}^{s,p}(\Omega)}\setminus\{0\}}\frac{[u]_{s,p}^{p}}{\left(\int_{\Omega}\frac{\lvert u(x)\rvert^{q}}{\lvert x\rvert^{m}}\,{\rm d}x\right)^{\frac{p}{q}}}\text{ and }S_{\alpha,\beta}:=\inf_{u,v\in W_{0}^{s,p}(\Omega)\setminus\{0\}}\frac{\lVert(u,v)\rVert^{p}_{{\mathcal{W}}}}{\left(\int_{\Omega}\frac{\lvert u\rvert^{\alpha}\lvert v\rvert^{\beta}}{\lvert x\rvert^{m}}\,{\rm d}x\right)^{\frac{p}{q}}}.$
Observe that, if $u$ minimizes $S_{\alpha+\beta}$, then so is $\lvert u\rvert$, since $\lvert u\rvert]_{s,p}\leq[u]_{s,p}$. So, without loss of generality, we assume that a minimizer of $S_{\alpha+\beta}$ is nonnegative. A similar fact holds for a minimizer of $S_{\alpha,\beta}$. If $(u,v)\in\mathcal{W}$ minimizes $S_{\alpha,\beta}$, then it satisfies
@@
$\left\{\begin{aligned} &(-\Delta_{p})^{s}u=\eta\lvert u\rvert^{r-2}u+\gamma\frac{\lvert u\rvert^{p^{*}_{s}(m)-2}u}{\lvert x\rvert^{m}}\;\;\text{in}\;\Omega,\\ &u=v=0\,\mbox{ in }\mathbb{R}^{d}\setminus\Omega.\end{aligned}\right$. (1.2)
@@
$\left\{\begin{aligned} &(-\Delta_{p})^{s}u=\eta_{1}\lvert u\rvert^{q-2}u+\frac{\alpha}{p_{s}^{*}}|u|^{\alpha-2}u|v|^{\beta}\;\;\text{in}\;\Omega,\\ &(-\Delta_{p})^{s}v=\eta_{2}\lvert v\rvert^{q-2}v+\frac{\beta}{p^{*}_{s}}|v|^{\beta-2}v|u|^{\alpha}\;\;\text{in}\;\Omega,\\ &u=v=0\;\;\text{in}\;\mathbb{R}^{d}\setminus\Omega,\end{aligned}\right$. (1.3)
@@
In this paper, we extend the result of [14, Theorem 1.2] in the context of the system. Specifically, we prove the existence of a nontrivial solution with the least energy for ($\mathcal{S}$). Moreover, we prove that the nontrivial solution is indeed non semi-trivial. To prove the non semi-triviality, we note that the approach used in [13], which relies on the presence of a concave nonlinearity, does not apply to our setting. Instead, we employ an alternative approach based on the properties of the Nehari manifold.
@@
[…]
@@
[…]
@@
$\Delta_{p}u=\mu u^{q-1}+u^{p^{*}-1}\text{ in }\Omega,\quad 2\leq p\leq q<p^{*}=\frac{dp}{d-p}.$
@@
$\Delta_{p}u+(-\Delta_{p})^{s}u=\lambda|u|^{q-2}u+|u|^{p^{*}-2}u\text{ in }\Omega,\;u=0\text{ in }\mathbb{R}^{d}\setminus\Omega,\quad 2\leq p<q<p^{*}=\frac{dp}{d-p},$
@@
Let $\Omega$ be a bounded open set containing zero and $m=0$. Let $2\leq p\leq r<p_{s}^{*}$, and $\alpha,\beta\geq 2$ be such that $\alpha+\beta=p^{*}_{s}$. Let $\gamma>0$, $r$ be as in Theorem 1.4 and $c_{0}$ be as in (1.5). Then there exists $\eta_{*}>0$ satisfying
@@
    For $\in(1,\infty), a:=a$$^{\prime}:=\frac{a}{a-1}$ is the conjugate exponent of $a$.
@@ 2. Existence and Non-existence of ground state solutions
$\int_{\Omega}\frac{\lvert u(x)\rvert^{\alpha}\lvert v(x)\rvert^{\beta}}{\lvert x\rvert^{m}}\,{\rm d}x=\int_{\Omega}\frac{\lvert u(x)\rvert^{\alpha}}{\lvert x\rvert^{\frac{m\alpha}{q}}}\frac{\lvert v(x)\rvert^{\beta}}{\lvert x\rvert^{\frac{m\beta}{q}}}\,{\rm d}x$
@@
Let $0\leq m\leq sp$, $p\leq q<p^{*}_{s}(m)$ and $\alpha,\beta>1$ be such that $\alpha+\beta=q$. Let $\{(u_{n},v_{n})\}$ weakly converges to (u,v){(u,v)} in $\mathcal{W}$. Then
@@
    Let ${\mathcal{M}}=\left\{(u,v)\in{\mathcal{W}}:G(u,v)=1\right\}$. Then $S_{\alpha,\beta}=\inf\left\{\lVert(u,v)\rVert^{p}_{{\mathcal{W}}}:(u,v)\in{\mathcal{M}}\right\}$ admits a positive minimizer in $\mathcal{M}$}.
@@
Similarly, we also get $\int_{\Omega}\frac{\lvert v_{n}-v\rvert^{q}}{|x|^{m}}\,{\rm d}x\rightarrow 0$ as $n\rightarrow\infty$. Therefore, taking $n\rightarrow\infty$ in (2.7) and (2.8), we conclude that $G$ is weakly continuous.
@@
Let $0\leq m\leq sp$, $p\leq q\leq p^{*}_{s}(m)$ and $\alpha,\beta>1$ be such that $\alpha+\beta=q$. Let $w_{0}$ be a minimizer of $S_{\alpha+\beta}$ and $(u_{0},v_{0})$ be a minimizer of $S_{\alpha,\beta}$. Then for every $B,C>0$ with $\frac{B}{C}=(\alpha\beta^{-1})^{\frac{1}{p}}$,
@@
$\text{for }\tilde{t}=\left(\frac{X}{Y}\right)^{\frac{1}{q-p}},\;E(\tilde{t}u,\tilde{t}v)=\max_{t>0}E(tu,tv).$
@@
Let $m=0$, i.e. $\alpha+\beta=p^{*}_{s}$. Then $S_{\alpha,\beta}$ has no minimizer on any open set $\Omega\subsetneq\mathbb{R}^{d}$.
@@
Observe that $\|(\cdot,\cdot)\|_{{\mathcal{W}}}$ and $\left(\int_{\mathbb{R}^{d}}|\cdot|^{\alpha}|\cdot|^{\beta}\,{\rm d}x\right)^{\frac{1}{p_{s}^{*}}}$ are invariant under translation and scaling
@@
Arguing as in [32, Section I.4.5, pg. 42], $S_{\alpha,\beta}$ is independent of $\Omega$ in $\mathbb{R}^{d}$. Suppose $(u_{0},v_{0})$ is a minimizer of $S_{\alpha,\beta}$ on any open set $\Omega$ in $\mathbb{R}^{d}$. Then the pair of functions $(\tilde{u},\tilde{v})\in\mathcal{W}(\mathbb{R}^{d})$ with $\tilde{u}=u_{0},\tilde{v}=v_{0}$ in $\Omega$, and $\tilde{u},\tilde{v}=0$ in $\mathbb{R}^{d}\setminus\Omega$, minimizes $S_{\alpha,\beta}$ over $\mathbb{R}^{d}$. Moreover, up to the normalized pair (as in Remark 1.1), they satisfy
$\left\{\begin{aligned} &(-\Delta_{p})^{s}\tilde{u}=\frac{\alpha}{p^{*}_{s}}|\tilde{u}|^{\alpha-2}\tilde{u}|\tilde{v}|^{\beta}\;\;\text{in}\;\mathbb{R}^{d},\\ &(-\Delta_{p})^{s}\tilde{v}=\frac{\beta}{p^{*}_{s}}|\tilde{v}|^{\beta-2}\tilde{v}|\tilde{u}|^{\alpha}\;\;\text{in}\;\mathbb{R}^{d},\\ &\tilde{u},\tilde{v}\geq 0\;\;\text{in}\;\mathbb{R}^{d}.\end{aligned}\right$. (2.11)
@@
Next, we recall the Morrey space.
Definition 2.7 (Morrey space).
For $r\in[1,\infty)$ and $\gamma\in[0,d]$, the homogeneous Morrey space $\mathcal{L}^{r,\gamma}(\mathbb{R}^{d})$ is defined as
$\mathcal{L}^{r,\gamma}(\mathbb{R}^{d})\coloneqq\left\{u:\mathbb{R}^{d}\to{\mathbb{R}}\text{ measurable }:\sup_{x\in\mathbb{R}^{d},\,{R>0}}R^{\gamma}\fint_{B(x,R)}|u(y)|^{r}\,{\rm d}y<\infty\right\},$
with norm
$\lVert u\rVert_{\mathcal{L}^{r,\gamma}(\mathbb{R}^{d})}^{r}\coloneqq\sup_{x\in\mathbb{R}^{d},\,{R>0}}R^{\gamma}\fint_{B(x,R)}|u(y)|^{r}\,{\rm d}y.$
In view of [8, Proposition 2.15], and using the fact that $W_{0}^{s,p}(\Omega)\subset{\mathcal{D}}^{s,p}(\mathbb{R}^{d})$, we have the following embedding.
@@
Let $\frac{p}{p^{*}_{s}}\leq\theta<1$ and $p\leq r\leq p^{*}_{s}$. Then
$\lVert(u,v)\rVert_{L^{p^{*}_{s}}(\mathbb{R}^{d})\times L^{p^{*}_{s}}(\mathbb{R}^{d})}$ $\leq C(d,p,s)\lVert(u,v)\rVert_{{\mathcal{W}}}^{\theta}$
$\lVert(u,v)\rVert_{\mathcal{L}^{r,\frac{r(d-sp)}{p}}(\mathbb{R}^{d})\times\mathcal{L}^{r,\frac{r(d-sp)}{p}}(\mathbb{R}^{d})}^{1-\theta},\;\forall\,(u,v)\in{\mathcal{W}}.$ (2.12)
Proposition 2.9.
Let $m=0$, i.e. $\alpha+\beta=p^{*}_{s}$. Let $\{(u_{n},v_{n})\}$ be a sequence in $\mathcal{W}$} such that
@@
[…]
@@
Therefore, $(u_{n},v_{n})\not\rightarrow 0$ in $L^{p^{*}_{s}}(\mathbb{R}^{d})\times L^{p^{*}_{s}}(\mathbb{R}^{d})$, and
$\inf_{n\in\mathbb{N}}\lVert(u_{n},v_{n})\rVert_{L^{p^{*}_{s}}(\mathbb{R}^{d})\times L^{p^{*}_{s}}(\mathbb{R}^{d})}^{p^{*}_{s}}\geq 1.$
Hence, for $\frac{p}{p^{*}}\leq\theta<1$, the interpolation given in (2.8) yields
$1\leq C\lVert(u_{n},v_{n})\rVert_{{\mathcal{W}}}^{\theta}\lVert(u_{n},v_{n})\rVert_{\mathcal{L}^{p,d-sp}(\mathbb{R}^{d})\times\mathcal{L}^{p,d-sp}(\mathbb{R}^{d})}^{1-\theta}\leq C_{1}\lVert(u_{n},v_{n})\rVert_{\mathcal{L}^{p,d-sp}(\mathbb{R}^{d})\times\mathcal{L}^{p,d-sp}(\mathbb{R}^{d})}^{1-\theta},$
which implies $\lVert(u_{n},v_{n})\rVert_{\mathcal{L}^{p,d-sp}(\mathbb{R}^{d})\times\mathcal{L}^{p,d-sp}(\mathbb{R}^{d})}\geq C_{2}$ for some $C_{2}>0$, i.e.,
$\sup_{x\in\mathbb{R}^{d},\,R>0}R^{d-sp}\fint_{B(x,R)}\left(|u_{n}(y)|^{p}+|v_{n}(y)|^{p}\right)\,{\rm d}y\geq C_{2}.$
Hence for every $n\in{\mathbb{N}}$, there exist $y_{n}\in\mathbb{R}^{d}$ and $r_{n}\in(0,\infty)$ such that
$r_{n}^{d-sp}\fint_{B(y_{n},r_{n})}\left(|u_{n}(y)|^{p}+|v_{n}(y)|^{p}\right)\,{\rm d}y\geq\lVert(u_{n},v_{n})\rVert_{\mathcal{L}^{p,d-sp}(\mathbb{R}^{d})\times\mathcal{L}^{p,d-sp}(\mathbb{R}^{d})}^{p}-\frac{C_{2}}{2n}>C_{3},$ (2.13)
for some $C_{3}>0$. By the change of variable and using the compact embedding of ${\mathcal{W}}\hookrightarrow L^{p}(\Omega)\times L^{p}(\Omega)$, we get
$\omega_{d}C_{3}\leq r_{n}^{-sp}\int_{B(y_{n},r_{n})}\left(|u_{n}(y)|^{p}+|v_{n}(y)|^{p}\right)\,{\rm d}y$ $=\int_{B(0,1)}\left(|\tilde{u}_{n}(y)|^{p}+|\tilde{v}_{n}(y)|^{p}\right)\,{\rm d}y$
$\longrightarrow\int_{B(0,1)}\left(|\tilde{u}(y)|^{p}+|\tilde{v}(y)|^{p}\right)\,{\rm d}y,\text{ as }n\rightarrow\infty.$
Therefore, $\tilde{u},\tilde{v}\neq 0$.
By Brezis Lieb lemma,
@@
$1$ $=\lim_{n\to\infty}\int_{{\mathbb{R}}^{d}}|\tilde{u}_{n}|^{\alpha}|\tilde{v}_{n}|^{\beta}\,{\rm d}x=\lim_{n\to\infty}\int_{{\mathbb{R}}^{d}}|\tilde{u}_{n}-\tilde{u}|^{\alpha}|\tilde{v}_{n}-\tilde{v}|^{\beta}\,{\rm d}x+\int_{{\mathbb{R}}^{d}}|\tilde{u}|^{\alpha}|\tilde{v}|^{\beta}\,{\rm d}x.$ (2.15)